OpenStudy (anonymous):

Use the five steps for problem solving to answer the following question. Please show all of your work. Ace Appliances charges a $45 fee plus $75 per hour for servicing refrigerators. Appliance Assessors charges a $60 fee plus $65 per hour. How many repair hours would make Appliance Assessors less expensive than Ace Appliances?

OpenStudy (anonymous):

Let t equal the number of hours where the two charge policies are equal to each other in cost. Solve the following equation for t: 45 + 75 t = 60 + 65 t t = 3/2 So t has to be greater than 3/2 repair hours to begin saving money using Appliance Assessors for servicing refrigerators.